A man is 20 years older than his son,the sum of their ages are 50. How old are they?
Answer by algebrahouse.com(1659)

x = son's age
x + 20 = man's age {he is 20 years older than his son}

x + x + 20 = 50 {sum of ages is 50}
2x + 20 = 50 {combined like terms}
2x = 30 {subtracted 20 from each side}
x = 15 {divided each side by 2}
x + 20 = 35 {substituted 15, in for x, into x + 20}

son is 15
man is 35

For more help from me, visit: www.algebrahouse.com